A hot-air balloonist, rising vertically with a constant speed of 5.00 m/s, releases a sandbag at the instant the balloon is 60.0 m above the ground. After it is released, the sandbag encounters no appreciable air resistance.

a) Find the position and velocity of the sandbag at 0.200 s and 0.800 s after its release.

b) How much time does it take for the sandbag to strike the ground after it is released?
Answer #1

(a)

Position of sandbag 0.2 s after its release,

y = yo + vo*t + (1/2)at^2

y = 60 + 5*0.2 - (1/2)*9.8*0.22

y = 60.804 m

Position 0.8 s after its release,

y = yo + vo*t + (1/2)at^2

y = 60 + 5*0.8 - (1/2)*9.8*0.82

y = 60.86 m

Velocity of sandbag after 0.2 s,

v = vo + at

v = 5 - 9.8*0.20

v = 3.04 m/s

Velocity after 0.8 s,

v = 5 - 9.8*0.80

v = -2.84 m/s

(b)

Consider upward direction positive and downward direction negative .

From kinematic equation,

h = vo*t + (1/2)at^2

60 = -5*t + (1/2)*9.8*t^2

4.9t^2 - 5t - 60 = 0

By solving above quadratic equation,

t = 4.046 s